# Concept-Based Interpretable Reinforcement Learning with Limited to No Human Labels

### v1

@@ 1. RQ 1
    Does LICORICE enable both high concept accuracy and high environment reward?
@@ 2. RQ 2
    To what extent does LICORICE enable a reduction in human labeling effort without substantially impacting performance?
@@ 3. RQ 3
    Can LICORICE be used alongside powerful vision-language models to further alleviate the labeling burden?
4.1 Environments
We evaluate our approach on three environments. For each environment, we define an interpretable concept set, where we define concepts to describe properties of objects of interest.
PixelCartPole
In PixelCartPole-v1 [31], the states are images, and the concepts are the original continuous features in the standard CartPole environment: the cart position, the cart velocity, the pole angle, and the pole angular velocity. To obtain the mapping from images to concepts, we use a fixed window of the four most recent images for the temporal concepts, such as the cart velocity. We also include the last action as input to ensure that the information is sufficient to infer concept values. This domain is deceptively difficult due to the temporal concepts.
DoorKey
In DoorKey-7x7 [5], the agent operates in a 5x5 grid [^1] In both of the Minigrid environments, the actual usable area is smaller than what shows in its name, as the outermost layer is a boundary. to pick up an item to unlock the door, then reach the green goal square. The states are fully observable images. We define 6 groups of concepts with 12 concept variables in total: 1) the agent’s position, 2) the agent’s direction, 3) the key’s position or a fixed position (0, 0) outside the grid if the key has been picked up, 4) the door’s position, 5) whether the door is open, and 6) whether the agent can move along each of four directions, i.e., the corresponding cell is empty. Each group contains one or more concept variables. For example, position groups contain x and y coordinates.
DynamicObstacles
In DynamicObstacles-5x5 [5], the agent operates in a 3x3 grid to avoid two moving obstacles while navigating to the goal. Colliding with the dynamic obstacles yields a large penalty, so the agent must correctly learn concepts to safely reach the goal. The states are fully observable images. We define 5 groups of concepts with 11 concept variables in total: 1) the agent’s position, 2) the agent’s direction, 3) the first obstacle’s position, 4) the second obstacle’s position, and 5) whether the agent can move along each of four directions, i.e., the corresponding cell is empty. Each group contains one or more concept variables. For example, position groups contain x and y coordinates.
4.2 Experiment Details
@@
In each experiment, we run each algorithm 5 times, each with a random seed. More implementation details and hyperparameters are in Section A.2.
Architecture
For the backbone model and algorithm, we use PPO (implementation provided by Stable Baselines 3 [23]). We preserve most of the network architecture, but we add a concept layer in the policy network. Specifically, after passing the image through the feature extractor, which consists of 3 CNN layers, we introduce a linear layer to map to the concept layer and obtain predicted concept values. Each predicted concept is represented as a real value for the regression case and one categorical value derived from a classification head for the classification case. This portion of the network corresponds to $g$. For $f$, we utilize an MLP extractor with 2 fully connected layers, each with 64 neurons and a Tanh activation function. We share the default CNN feature extractor between policy and value functions, as we found this to be beneficial in preliminary experiments.
Concept Representation
We model concept learning for PixelCartPole-v1 as a regression problem (minimizing mean squared error) as the concepts are real-valued, and concept learning for DoorKey-7x7 and DynamicObstacles-5x5 as classification problems as the concepts are categorical. The observation images have resolutions of $(240,160)$ for PixelCartPole-v1, and $(160,160)$ for DoorKey-7x7 and DynamicObstacles-5x5. These dimensions maintain a short side of 160 pixels while preserving the aspect ratio of the original rendered images, as we observe lower resolutions may impact performance.
Performance Metrics
@@
To obtain an upper bound on the reward for each environment, we directly use ground-truth concept labels to learn a policy, leading to a reward of $500$, $0.97$, and $0.91$ for the three environments respectively. In the following sections, we report the reward as a ratio of this upper bound. Percentages (or ratios) make sense here since the minimum reasonable reward is 0 in all environments: in PixelCartPole-v1 and DoorKey-7x7, all rewards are nonnegative; in DynamicObstacles-5x5, a random policy would have negative reward due to collisions, but the agent can ensure nonnegative reward by simply staying in place. We additionally report the concept error (MSE for regression; $1$ accuracy for classification). All the reported numbers are calculated during the testing stage, where we evaluate the models on 100 episodes.
@@ 4.3 Results
PixelCartPole-v1 DoorKey-7x7 DynamicObstacles-5x5
$c$ Labels Algorithm $R$uparrow $c$ MSE $\downarrow$ $R$uparrow $c$ Error $\downarrow$ $R$uparrow $c$ Error $\downarrow$
$B$ Sequential-Q 0.232 4720.2324720.232\,4720.232 472 0.102 570.102570.102\,570.102 57 0.462 5660.4625660.462\,5660.462 566 0.462 3270.4623270.462\,3270.462 327 0.788 3480.7883480.788\,3480.788 348 0.014 1360.0141360.014\,1360.014 136
Disagreement-Q 0.363 6240.3636240.363\,6240.363 624 0.097 4570.0974570.097\,4570.097 457 0.759 4640.7594640.759\,4640.759 464 0.392 3940.3923940.392\,3940.392 394 0.992 530.992530.992\,530.992 53 0.000 7790.0007790.000\,7790.000 779
Random-Q 0.337 720.337720.337\,720.337 72 0.065 6130.0656130.065\,6130.065 613 0.886 5670.8865670.886\,5670.886 567 0.269 730.269730.269\,730.269 73 0.981 0960.9810960.981\,0960.981 096 0.022 4280.0224280.022\,4280.022 428
LICORICE $1.0$ 0.024 7970.0247970.024\,7970.024 797 0.990 2970.9902970.990\,2970.990 297 0.047 5040.0475040.047\,5040.047 504 0.990 6070.9906070.990\,6070.990 607 0.000 1930.0001930.000\,1930.000 193
$\infty$ CPM  [33] $1.0$ 0.011 5450.0115450.011\,5450.011 545 1.001 0241.0010241.001\,0241.001 024 $0.0$ 0.969 5770.9695770.969\,5770.969 577 $0.0$
Table 1: Evaluation of the reward $R$ and concept error achieved by all methods in all environments. The reward is reported as the fraction of the reward upper bound. Unless otherwise stated, concept error is 1 - accuracy. Methods in the top section are given a budget of $B=[500,300,300]$, respectively; CPM is given an unlimited budget (in practice, it uses 4M, 4M, 1M concept labels respectively). For each column, the top-performing method with a limited budget is in bold. If there are several methods with overlapping standard deviations, we bold all of them. Full results with standard deviation are in Table 6, Appendix B.
4.3.1 Balancing Concept Performance and Environment Reward
We first validate that LICORICE can achieve high reward in all test environments while accurately identifying concepts. We first compare against baselines with a fixed budget of $B=[500,300,300]$ queries for PixelCartPole-v1, DoorKey-7x7, and DynamicObstacles-5x5, respectively, as depicted in the first section of Table 1. We choose these budgets by starting from $500$ and then decreasing by units of $100$ until we find that LICORICE can no longer achieve $99\%$ of the reward upper bound if we further decrease the budget. We treat the minimum budget under which LICORICE can achieve $99\%$ of the reward upper bound as the maximum budget considered for that environment.
Comparison with Budget-Constrained Baselines
To our knowledge, no previous algorithms exist that seek to minimize the number of train-time labels for concept-based RL training, so we implement our own baselines. In Sequential-Q, the agent spends all of $B$ queries on the first $B$ states it encounters during the initial policy rollout. In Disagreement-Q, the agent similarly spends its budget at the beginning of its learning process; however, it uses active learning with the same acquisition function as LICORICE to strategically choose the training data. In Random-Q, the agent receives $B$ concept labels at random points in the training process using a probability to decide whether to query for a concept for each state. We show the results in the first section of Table 1. In the three environments, LICORICE performs similarly to or better than all baselines in terms of both reward and concept error. The main performance differences can be seen for PixelCartPole-v1 and DoorKey-7x7, where LICORICE achieves $100\%$ and $99\%$ of the optimal reward, respectively, while the second best algorithm achieves $36\%$ and $89\%$, respectively. Random-Q, Disagreement-Q, and LICORICE perform similarly on DynamicObstacles-5x5, indicating that this environment is relatively simple and may not benefit from a more advanced strategy. Its simplicity may be due to the lack of concept distribution shift that is present in the other two environments. For example, in PixelCartPole-v1, the policy must be further refined to improve its estimate of the more on-policy concepts. In the initial training stages, when the policy is more random, the agent needs to estimate a wide range of possible values for the pole angle, as the pole starts upright and frequently falls. As the agent’s policy improves, it requires more precise estimates of the pole angle primarily around the vertical position.
Comparison with Budget-Unconstrained Baseline
To better understand the performance of our approach, we compare it against existing work that is not constrained by concept label budgets. Specifically, we implement Concept Policy Model (CPM) from previous work in multi-agent RL [33] but for the single-agent setting. This approach jointly trains the concept bottleneck and the policy, assuming unlimited access to concept labels. It represents an upper bound on concept accuracy, as the agent receives continuous concept feedback throughout learning. Surprisingly, LICORICE outperforms CPM in PixelCartPole-v1and has similar performance to CPM in the other two environments in this unfair comparison. We emphasize that CPM is given an unlimited budget, and in fact, it uses over 1M concept labels for each environment, whereas LICORICE uses $500$ or fewer. We therefore answer RQ 1 affirmatively: compared to baselines, LICORICE demonstrates both high concept accuracy (low concept error) and high reward on our three test environments, all while using substantially fewer concept labels than existing algorithms.
PixelCartPole-v1 DoorKey-7x7 DynamicObstacles-5x5
$B$ 300 400 500 100 200 300 100 200 300
$R$ 0.276 3440.2763440.276\,3440.276 344 0.766 2880.7662880.766\,2880.766 288 $1.0$ 0.765 3910.7653910.765\,3910.765 391 0.934 5320.9345320.934\,5320.934 532 0.990 2970.9902970.990\,2970.990 297 0.957 4390.9574390.957\,4390.957 439 0.979 8190.9798190.979\,8190.979 819 0.990 6070.9906070.990\,6070.990 607
$c$ Error 0.153 5730.1535730.153\,5730.153 573 0.052 4580.0524580.052\,4580.052 458 0.024 7970.0247970.024\,7970.024 797 0.260 4040.2604040.260\,4040.260 404 0.086 6350.0866350.086\,6350.086 635 0.047 5040.0475040.047\,5040.047 504 0.040 7940.0407940.040\,7940.040 794 0.007 8340.0078340.007\,8340.007 834 0.000 1930.0001930.000\,1930.000 193
Table 2: Performance of LICORICE on all environments for varying maximum budgets. The reward is reported as the fraction of the reward upper bound. For PixelCartPole-v1, $c$ error is MSE; for the other environments, $c$ error is 1 - accuracy. For all environments, the reward increases and the error decreases as we increase the maximum budget. Full results with standard deviation are in Table 7Appendix B.
4.3.2 Budget Allocation Effectiveness
To understand the extent to which LICORICE enables a reduction in human labeling effort without substantially impacting performance, we conduct experiments across three different environments with varying concept labeling budgets. We start with the baseline budget $B$ from our previous experiments and test two additional budgets, decreasing by steps of 100. The only component of our algorithm that we vary is the number of iterations $M$. For each budget and environment, we report a single performance value that maximizes the sum of the relative reward and inverse concept error, both scaled to $[0,1]$. The results of this experiment are shown in Table 2.
Our findings reveal that the impact of budget reductions (human labeling effort) varies across environments. In general, we see an increase in concept error and a decrease in reward for all environments with decreasing budget. However, both DynamicObstacles-5x5 and DoorKey-7x7 show more resilience to budget reductions, while PixelCartPole-v1 exhibits higher sensitivity. Specifically, we find that even with $B=100$ for DynamicObstacles-5x5, LICORICE can still achieve $96\%$ of the optimal reward. With $B=200$, LICORICE can still achieve $93\%$ of the optimal reward on DoorKey-7x7. In contrast, we see large decreases in reward from the budget reductions for PixelCartPole-v1. We therefore provide a more nuanced answer to RQ 2: LICORICE can indeed enable reductions in human labeling effort without substantially impacting performance, but the extent of this reduction is environment-dependent. In simpler environments, human effort can be dramatically reduced while maintaining near-optimal performance. However, in more complex environments like PixelCartPole-v1, the trade-off between reduced human effort and maintained performance becomes more pronounced. This finding aligns with the intuition that different environments have varying levels of complexity and would require different levels of human input to accurately capture all relevant concepts and their relationships.
Figure 1: Training curves of reward and concept error for PixelCartPole-v1. Shaded values correspond to 95% CI, calculated using 1000 bootstrap samples. All curves in Figure 5.
4.3.3 Integration with Vision-Language Models
@@
We now seek to answer the question of whether VLMs can successfully provide concept labels in lieu of a human annotator within our LICORICE framework. For consistency, we use the same hyperparameters for LICORICE as in Section 4.3.2. Because using VLMs incurs costs and users requiring interpretable policies for their environments may still face budget constraints, we operate within the same budget-constrained setting as described in Section 4.3.2. Here, the reduction in the number of “human” labels required translates to cost savings.
We show the results in Table 3. We find that GPT-4o can indeed serve as an annotator for some but not all environments. In DoorKey-7x7 and DynamicObstacles-5x5, LICORICE with GPT-4o labels achieves 83% and 88% of the optimal reward, respectively. The concept error rate is comparable to GPT-4o’s labeling error, both evaluated on the same rollout observations, with sampling applied to the latter due to cost constraints. Additionally, the concept error gap between our trained concept model and GPT-4o almost diminishes with more labels. Taken together, these results suggest that LICORICE can work reasonably well with GPT-4o labels.
In PixelCartPole-v1, however, there are challenges for GPT-4o in providing accurate labels due to the continuous nature of the concepts and the lack of necessary knowledge of physical rules and quantities in this specific environment. As a result, LICORICE with GPT-4o labels achieves generally low reward across varying budgets. Interestingly, the concept error increases with more budget, likely due to the additional labeling noise introduced to the concept network. This suggests that while the VLM can handle labeling tasks to some extent, its performance in PixelCartPole-v1 is limited by the complexity of accurately labeling concepts. We also posit the mean reward drop in DynamicObstacles-5x5 at 300 budget is due to the randomness introduced by GPT-4o; however, we note that the range of standard deviation overlaps with the other two budgets. We therefore answer RQ 3 with cautious optimism: there are indeed cases in which LICORICE could be used alongside VLMs. Generally speaking, a more human-intuitive and less complex environment is more likely to work well with VLMs by enabling clear and detailed labeling instructions in the prompt.
PixelCartPole-v1 DoorKey-7x7 DynamicObstacles-5x5
$B$ 300 400 500 100 200 300 100 200 300
$R$ (LICORICE GPT-4o) 0.064 7120.0647120.064\,7120.064 712 0.066 2160.0662160.066\,2160.066 216 0.064 940.064940.064\,940.064 94 0.711 4030.7114030.711\,4030.711 403 0.735 0890.7350890.735\,0890.735 089 0.830 1130.8301130.830\,1130.830 113 0.268 7110.2687110.268\,7110.268 711 0.926 0950.9260950.926\,0950.926 095 0.875 0150.8750150.875\,0150.875 015
$c$ Error (LICORICE GPT-4o) 0.170 0680.1700680.170\,0680.170 068 0.246 0870.2460870.246\,0870.246 087 0.246 895 7330.2468957330.246\,895\,7330.246 895 733 0.450 3260.4503260.450\,3260.450 326 0.329 3480.3293480.329\,3480.329 348 0.312 2880.3122880.312\,2880.312 288 0.200 7110.2007110.200\,7110.200 711 0.133 8850.1338850.133\,8850.133 885 0.132 220.132220.132\,220.132 22
$c$ Error (GPT-4o) 0.193 750.193750.193\,750.193 75 0.224 6830.2246830.224\,6830.224 683 0.241 3650.2413650.241\,3650.241 365 $0.313$ 0.313 6670.3136670.313\,6670.313 667 $0.304$ 0.159 6360.1596360.159\,6360.159 636 0.123 6360.1236360.123\,6360.123 636 0.127 6360.1276360.127\,6360.127 636
Table 3: Performance of LICORICE with GPT-4o integrated into the loop for all environments across different budgets. We also report the concept labeling error on a random sample of 50 observations from the same rollout set used for LICORICE+GPT-4o, due to API cost constraints. Full results with standard deviation are in Table 8, Appendix B.
PixelCartPole-v1 DoorKey-7x7 DynamicObstacles-5x5
Algorithm $R$uparrow $c$ MSE $\downarrow$ $R$uparrow $c$ Error $\downarrow$ $R$uparrow $c$ Error $\downarrow$
LICORICE-IT 0.533 8960.5338960.533\,8960.533 896 0.079 6220.0796220.079\,6220.079 622 0.992 4120.9924120.992\,4120.992 412 0.089 5350.0895350.089\,5350.089 535 0.996 8890.9968890.996\,8890.996 889 0.001 2020.0012020.001\,2020.001 202
LICORICE-DE $0.9669$ 0.031 1060.0311060.031\,1060.031 106 0.783 1130.7831130.783\,1130.783 113 0.200 0220.2000220.200\,0220.200 022 0.980 5370.9805370.980\,5370.980 537 0.000 1140.0001140.000\,1140.000 114
LICORICE-AC 0.910 7840.9107840.910\,7840.910 784 0.022 3140.0223140.022\,3140.022 314 0.815 7550.8157550.815\,7550.815 755 0.161 8420.1618420.161\,8420.161 842 0.583 0530.5830530.583\,0530.583 053 0.002 2910.0022910.002\,2910.002 291
LICORICE $1.0$ 0.024 7970.0247970.024\,7970.024 797 0.990 2970.9902970.990\,2970.990 297 0.047 5040.0475040.047\,5040.047 504 0.990 6070.9906070.990\,6070.990 607 0.000 1930.0001930.000\,1930.000 193
@@
Table 4: Ablation study results for LICORICE in all environments. All of our components generally help achieve a better reward and lower concept error. Full results with standard deviation are in Table 9, Appendix B.
4.3.4 Ablation Study
Given the previous results, we now conduct ablations to confirm the effectiveness of our three main contributions: iterative training, decorrelation, and active learning. LICORICE-IT corresponds to LICORICE with only one iteration, LICORICE-DE corresponds to LICORICE without decorrelation, and LICORICE-AC corresponds to LICORICE without active learning (instead, it uses the entire unlabeled dataset for querying). We show the learning curves for PixelCartPole-v1 in Figure 1; the full set of curves is in Appendix B.
@@ 4.3.4 Ablation Study
Table 4 depicts the results of our ablation study on all environments. The bottom row corresponds to the upper bound on performance by LICORICE, with all components included. All of our contributions are critical to achieving both high reward and low concept error. However, the component that most contributes to the reward or concept performance differs depending on the environment. For example, compared with LICORICE, LICORICE-IT exhibits the largest reward gap for PixelCartPole-v1; however, LICORICE-AC yields the largest reward gap for DynamicObstacles-5x5, and LICORICE-DE yields the largest gap for DoorKey-7x7. We suspect that this difference is because the concepts in DynamicObstacles-5x5 are simple enough such that one iteration is sufficient for learning, meaning that the largest gains can be made by using active learning. In contrast, PixelCartPole-v1 requires further refinement of the policy to better estimate on-distribution concept values, so the largest gains can be made by leveraging multiple iterations.

### v2

@@ 1. RQ 1
    Under a limited concept annotation budget, does LICORICE enable both high concept accuracy and high environment reward?
@@ 2. RQ 2
    How effective are VLMs as automated concept annotators when used with LICORICE?
@@ 3. RQ 3
    How label efficient is LICORICE compared with other methods?
  4. RQ 4
    Does LICORICE support test-time concept interventions?
4.1 Experiment Setup
@@
In each experiment, we run each algorithm 5 times, each with a random seed. All algorithms use PPO (Schulman et al., 2017; Raffin et al., 2021) with a concept bottleneck. More implementation details and hyperparameters are in Section A.2.
Environments. We investigate these questions across five environments: PixelCartPole (Yang et al., 2021), DoorKey (Chevalier-Boisvert et al., 2023), DynamicObstacles (Chevalier-Boisvert et al., 2023), Boxing (Bellemare et al., 2013), and Pong (Bellemare et al., 2013). Each environment includes a distinct challenge and features a set of interpretable concepts describing key objects and properties. We summarize the concepts in Table 1, with more details in Section A.1. These environments are characterized by their dual representation: a complex image-based input and a symbolic concept-based representation. PixelCartPole, DoorKey, and DynamicObstacles are simpler because we can extract noiseless ground-truth concept labels from their source code. In contrast, Boxing, as implemented in OCAtari (Delfosse et al., 2023), uses reverse engineering to extract positions of important objects from the game’s RAM state. This extraction process introduces a small amount of noise. The concepts in Pong are derived from the VIPER paper (Bastani et al., 2018) that also uses reverse engineering.
Environment # Concepts Concept Type Binarized # Concepts
PixelCartPole 4 Continuous N/A
DoorKey 12 Discrete 46
DynamicObstacles 11 Discrete 30
Boxing 8 Discrete 1480
Pong 7 Discrete 1848
Table 1: Summary of environments, including their associated concepts. Counts for the binarized version of the concepts provided where applicable to illustrate the problem size.
Baselines. To our knowledge, there exist no previous algorithms that seek to minimize the number of concept labels for interpretable RL, so we implement three: Sequential-Q, Disagreement-Q, and Random-Q. In Sequential-Q, the agent spends $B$ queries on the first $B$ states encountered during the initial policy rollout. In Disagreement-Q, the agent also spends its budget at the beginning of training; however, it uses active learning with the same $\alpha(\cdot)$ as LICORICE to strategically choose the training data for annotation. In Random-Q, the agent receives $B$ concept labels at random points using a probability to decide whether to query in each state. As a budget-unconstrained baseline, we implement CPM from previous work in multi-agent RL (Zabounidis et al., 2023) for the single-agent setting. As mentioned in §2, CPM jointly trains $g$ and $f$, assuming unlimited access to concept labels, so it also represents an upper bound on concept accuracy.
VLM Details. For our VLM experiments, we use GPT-4o (gpt, ). During training, we query GPT-4o each time LICORICE requires a concept label. As an example, for the concept related to the position of an object, we prompt GPT-4o with instructions to report the coordinates (x y). More details about our prompts can be found in Section A.3.
@@
Performance Metrics. We compare our reward to an upper bound calculated by training PPO with ground-truth concept labels (PixelCartPole: $500$, DoorKey: $0.97$, DynamicObstacles: $0.91$, Boxing: $86.05$, Pong: $21$). In the first four environments, percentages (or ratios) make sense since the minimum reasonable reward is 0.222In PixelCartPole and DoorKey, all rewards are nonnegative; in DynamicObstacles, an agent can ensure nonnegative reward by simply staying in place; in Boxing, a random policy can get around 0 reward and all of the trained policies have positive reward. In Pong, a random policy has -21 reward so we first get the difference between rewards and -21 and then calculate the ratio. We additionally report the concept error (MSE for regression; $1\,-\,$accuracy for classification). In Boxing, following the practice of OCAtari, we consider a concept prediction correct if it is within 5 pixels of the ground truth label. In Pong, we simply use classification error. All reported numbers are calculated using 100 evaluation episodes.
@@ 4.2 Results
Figure 2: LICORICE generally achieves higher reward and lower concept error than all other budget-constrained algorithms (Sequential-Q, Disagreement-Q, Random-Q). The budgets for each environment are PixelCartpole: $500$, DoorKey: $300$, DynamicObstacles: $300$, Boxing: $3000$, Pong: $5000$. CPM has an unlimited budget (in practice, it uses 4M, 4M, 1M, 15M, 10M labels, respectively). Despite this extreme budget difference, LICORICE achieves comparable reward and concept error. Full results in Table 5, Appendix B.
RQ 1: Reward and concept accuracy under a limited annotation budget
LICORICE achieves similarly high reward and concept performance to the state-of-the-art budget-unconstrained baseline.
We first seek to understand how LICORICE performs compared to the state-of-the-art in concept-based RL, CPM, which is not constrained by concept label budgets. Surprisingly, LICORICE and CPM achieve nearly 100% of the maximum reward in all environments in this unfair comparison (Figure 2). They also achieve a similar concept error rate, only seeing a clear difference in error for the most complex environment, Pong. We emphasize that CPM is given an unlimited budget; in practice, it uses 1M-15M concept labels for all environments, which is around $2000$ to 13000×13000$\times$13000 × the budget used by LICORICE. In contrast, LICORICE operates under a strict budget constraint, with $3000$ labels for Boxing and $5000$ for Pong, and at most $500$ concept labels for the simpler environments (PixelCartPole: $500$, DoorKey: $300$, DynamicObstacles: $300$).
LICORICE generally achieves higher reward and lower concept error than budget-constrained baselines. We study all budget-constrained algorithms (LICORICE plus those in §4.1) under the same fixed concept labeling budget as above. Figure 2 shows that LICORICE outperforms all baselines in terms of reward and concept error on all but arguably the easiest environment, DynamicObstacles. In that environment, LICORICE performs similarly to the baselines. The greatest performance differences occur in PixelCartPole and DoorKey, where LICORICE achieves $100\%$ and $99\%$ of the maximum reward, while the second-best algorithm achieves $36\%$ and $89\%$, respectively. We therefore answer RQ 1 affirmatively: not only does LICORICE achieve the same high concept accuracy and high reward as the state-of-the-art in concept-based RL, it also performs on-par to budget-constrained baselines in one environment and outperforms them in the other four.
RQ 2: VLMs as concept annotators
@@
We now seek to answer the question of whether VLMs can successfully provide concept labels in lieu of a human annotator within our LICORICE framework. Because using VLMs incurs costs, we operate within the same budget-constrained setting as above. We present these results in Figure 2.
$c$ Error
Environment LICORICE+GPT-4o GPT-4o
PixelCartPole 0.246 8960.2468960.246\,8960.246 896 0.241 3650.2413650.241\,3650.241 365
DoorKey 0.312 2880.3122880.312\,2880.312 288 0.304 $0$.3040000.304\,$0$.304 000
DynamicObstacles 0.132 220.132220.132\,220.132 22 0.127 6360.1276360.127\,6360.127 636
Boxing 0.983 520.983520.983\,520.983 52 0.815 8330.8158330.815\,8330.815 833
Pong 0.728 870.728870.728\,870.728 87 0.874 130.874130.874\,130.874 13
Table 2: Concept error comparison. The concept error of LICORICE+GPT-4o matches that of GPT-4o alone in all environments except Boxing.
VLMs can serve as concept annotators for some environments. In DoorKey and DynamicObstacles, LICORICE with GPT-4o labels achieves 83% and 88% of the maximum reward, respectively. To assess the quality of VLM-generated labels, we compare the concept error rate of our trained model against GPT-4o’s labeling error, both evaluated on the same rollout observations, while GPT-4o evaluation is on a random subset of 50 observations, due to API cost constraints. Table 2 shows that the concept error of LICORICE is comparable to that of GPT-4o when GPT-4o labeling error is not high, which suggests that LICORICE can effectively utilize these concept labels. VLMs, particularly GPT-4o, could serve as concept annotators for certain concepts in certain environments.
VLMs struggle to provide accurate concepts in more complex environments. In PixelCartPole, Boxing, and Pong environments, however, GPT-4o struggles to provide accurate labels. In these environments, not only does LICORICE+GPT-4o achieve less than $25\%$ of the maximum reward, it also incurs a large concept error. The challenge with PixelCartPole is the continuous nature of the concepts and the lack of necessary knowledge of physical rules and quantities in this specific environment. Boxing is particularly challenging due to the large number of possible concept values that each of the 8 discrete concepts can take on (160 or 210 possible values). Similarly, concepts in Pong have hundreds of possible values and some physics quantities require multi-frame inference. We therefore answer RQ 2 with cautious optimism: there are indeed cases in which LICORICE could be used alongside VLMs, though not all yet. Generally speaking, a less complex environment is more likely to work well with VLMs by enabling clear and detailed labeling instructions in the prompt. However, VLM capabilities need improvement before they can fully alleviate the human annotation burden, especially in safety-critical settings.
RQ 3: Investigation of Budget Requirements for Performance
Given these results, we now investigate the minimum budget required for all environments to achieve 99% of the reward upper bound. To do so, we incrementally increase the budget for each environment starting from a reasonably small budget until LICORICE reaches 99% of the reward upper bound. In Pong, we further increase the budget until the concept error is below $0.25$. In addition to the baselines, we study LICORICE under perfect (human) annotation and noisy VLM labels in Figure 3.
LICORICE more rapidly achieves high reward compared to baselines. Across all environments, LICORICE is the most query-efficient. In three of the environments, it consistently achieves higher reward than baselines across all budget levels. In DynamicObstacles and Pong, LICORICE outperforms baselines at the smallest query budget, after which some baseline method achieves comparable reward. LICORICE is similarly performant with respect to concept error. Except for one budget setting for one environment, LICORICE consistently achieves lower concept error than the baselines.
For LICORICE+GPT-4o, more concept labels is not always better. In DoorKey, LICORICE+GPT-4o exhibits predictable behavior in terms of concept error: as the budget increases, the reward increases and the concept error decreases. Counterintuitively, for some environments, the concept error and reward fluctuate with more budget, likely due to the additional labeling noise introduced. We therefore provide a more nuanced answer to RQ 3: LICORICE is indeed label efficient across varying budgets, but the benefit is annotator-dependent.
Figure 3: Reward (top) and concept error (bottom) of all algorithms for different budgets. LICORICE more efficiently makes use of the varying budgets, achieving higher reward and lower concept error. Full results with standard deviation are in Tables 6, 7 and 8, Appendix B.
RQ 4: Interpretability Analysis: Test-Time Concept Interventions
A great property of concept-based networks is the ability for people to successfully intervene on the concepts to correct them. In RL, this intervention enables the inspection of how different concepts influence the immediate decisions of the agent.
Figure 4: Concept intervention results: LICORICE enables test-time concept intervention.
LICORICE enables test-time concept intervention. To validate that LICORICE supports test-time concept intervention, we simulate using a noisy concept model $\hat{g}$ with $f$ to study the impact of concept interventions on reward. For PixelCartPole, we add Gaussian noise to each predicted concept with a standard deviation of $0.2$. For the other three environments with discrete concepts, each concept label is randomly changed with probability $0.2$. Following previous work (Koh et al., 2020), we first intervene individually on concepts, setting them to ground truth, and sort the concepts in descending order of reward improvement. We then sequentially intervene on the concepts following this ordering, such that any previously intervened concepts remain intervened. According to Figure 4, using a noisy concept model significantly degrades reward. However, the performance increases with more interventions, meaning LICORICE supports test-time concept intervention, affirmatively answering RQ 4.
Figure 5: Test-time intervention examples, where intervening on a single concept corrects the action.
Concept interventions show how the policy decisions change. We now show how domain experts could interact with the model. Specifically, we simulate a counterfactual question: What if a concept value is incorrectly predicted? Figure 5 depicts two examples in DoorKey. Just intervening on the concept of the door being open or closed is sufficient to change the agent’s behavior, both highlighting the importance of this specific concept and the ability of a user to intervene on the concepts to interrogate and understand agent behavior.
Additional experiments
PixelCartPole DoorKey DynamicObstacles Boxing
Algorithm $R$uparrow $c$ MSE $\downarrow$ $R$uparrow $c$ Error $\downarrow$ $R$uparrow $c$ Error $\downarrow$ $R$uparrow $c$ Error $\downarrow$
LICORICE-IT 0.533 8960.5338960.533\,8960.533 896 0.079 6220.0796220.079\,6220.079 622 0.992 4120.9924120.992\,4120.992 412 0.089 5350.0895350.089\,5350.089 535 0.996 8890.9968890.996\,8890.996 889 0.001 2020.0012020.001\,2020.001 202 0.978 1520.9781520.978\,1520.978 152 0.135 6060.1356060.135\,6060.135 606
LICORICE-DE $0.9669$ 0.031 1060.0311060.031\,1060.031 106 0.783 1130.7831130.783\,1130.783 113 0.200 0220.2000220.200\,0220.200 022 0.980 5370.9805370.980\,5370.980 537 0.000 1140.0001140.000\,1140.000 114 1.001 0451.0010451.001\,0451.001 045 0.044 7890.0447890.044\,7890.044 789
LICORICE-AC 0.910 7840.9107840.910\,7840.910 784 0.022 3140.0223140.022\,3140.022 314 0.815 7550.8157550.815\,7550.815 755 0.161 8420.1618420.161\,8420.161 842 0.583 0530.5830530.583\,0530.583 053 0.002 2910.0022910.002\,2910.002 291 0.994 3050.9943050.994\,3050.994 305 0.087 9810.0879810.087\,9810.087 981
LICORICE $1.0$ 0.024 7970.0247970.024\,7970.024 797 0.990 2970.9902970.990\,2970.990 297 0.047 5040.0475040.047\,5040.047 504 0.990 6070.9906070.990\,6070.990 607 0.000 1930.0001930.000\,1930.000 193 1.004 9971.0049971.004\,9971.004 997 0.049 $0$.0490000.049\,$0$.049 000
@@
Table 3: Ablation study results for LICORICE. All components generally help achieve better reward and lower concept error. Full results with standard deviation are in Table 9, Appendix B.
@@
Ablation study: all components of LICORICE contribute positively to its performance. We now conduct an ablation study of our three main contributions: iterative training, decorrelation, and disagreement-based active learning. LICORICE-IT uses only one iteration, LICORICE-DE does not perform data decorrelation, and LICORICE-AC uses the entire unlabeled dataset for querying instead of a subset chosen by active learning. All components are critical to achieving both high reward and low concept error (Table 3; learning curves in Appendix B). However, the component that most contributes to the performance differs depending on the environment. For example, compared with LICORICE, LICORICE-IT exhibits the largest reward gap for PixelCartPole; however, LICORICE-AC yields the largest reward gap for DynamicObstacles, and LICORICE-DE yields the largest gap for DoorKey. This difference may occur because the concepts in DynamicObstacles are simple enough that one iteration is sufficient, so the largest gains can be made with active learning. In contrast, PixelCartPole requires further policy refinement to better estimate on-distribution concept values, so the largest gains can be made by leveraging multiple iterations.
LICORICE is robust to various hyperparameter values. LICORICE involves a few key hyperparameters, described in §3 and summarized here for convenience. In data decorrelation, $\tau$ governs the size of the unlabeled dataset collected by the current policy and $p$ controls the rate of acceptance of data points. Active learning involves a concept ensemble with $N$ models. We study the effect of the values of these hyperparameters on DynamicObstacles, finding that LICORICE achieves $96-99\%$ of the maximum reward across $2$ values for $N$, $3$ values for $p$, and $3$ values for $\tau$ (Section B.5).